# Quantum Signatures of Chaos from Free Probability

### v1

@@ 1 Introduction
Despite its significance across various domains, quantum chaos remains poorly understood. Several diagnostics have been proposed, including spectral statistics governed by random matrix theory BGS , ETH DAlessio:2015qtq , OTOCs Shenker:2014cwa , Krylov complexityParker:2018yvk , and spread complexity Balasubramanian:2022tpr . While these probes offer valuable insights, their interconnections are not well established. [^1] For previous work in this direction, see Cotler:2017jue , which relates averaged OTOCs to the spectral form factor, and rozenbaum2019universal , which discusses the spectral statistics of Lyapunovian operators. These topics are reviewed in Appendices D and E, respectively. It remains unclear whether an overarching theoretical framework exists to unify these different manifestations of quantum chaos or at least provide a systematic classification of them. In many respects, our current understanding of quantum chaos is largely phenomenological, relying on conjectures and heuristic arguments often inspired by semiclassical intuition. For instance, periodic orbit theory has been successfully employed to derive random matrix behavior from classical periodic orbits Berry1985 . However, this approach inherently assumes the presence of well-defined classical trajectories and thus breaks down in systems without a classical limit. Moreover, the mathematically rigorous framework for characterizing chaos – quantum ergodic theory (see Gesteau:2023rrx ; Ouseph:2023juq for recent reviews) – remains largely disconnected from the aforementioned diagnostics of quantum chaos. Developing a mathematically precise and conceptually unified theory of quantum chaos is, therefore, a crucial open challenge, one that could provide a systematic framework for addressing many outstanding questions in the field.
@@
In this work, we propose a quantum ergodic theory based on free probability theory mingo2017free . This approach is motivated by the fact that any quantum dynamical system can be viewed as a non-commutative probability space equipped with a time evolution map. Since free probability theory studies precisely such non-commutative probability spaces, it is natural to expect its relevance in quantum-mechanical systems. By analogy with classical ergodic theory, chaos can be characterized by how much operators in the future, $B(t)$, become statistically independent of operators in the past, $A(0)$. In free probability theory, a fundamental notion of statistical independence is freeness, meaning that all mixed free cumulants involving the corresponding random variables vanish. In fact, various manifestations of quantum chaos, such as the ETH and the vanishing of OTOCs, can be understood as different avatars of freeness Jindal:2024zcg . Strictly speaking, freeness only emerges in the thermodynamic limit, where the rank $\mathcal{D}$ of the operators in a quantum many-body system tends to infinity. In this context, one speaks of asymptotic freeness, indicating that freeness holds as $\mathcal{D}\to\infty$. At finite $\mathcal{D}$, a hallmark of asymptotic freeness is that the spectrum of combinations of asymptotically free operators follows the predictions of free probability theory Chen:2024zfj . Another significant application of free probability theory in the context of quantum chaos is the development of generalized versions of ETH that aim to explain the behavior of OTOCs Pappalardi:2022aaz .
@@
This work is organized as follows. In Section 2, we review classical ergodic theory. In Section 3, we discuss quantum ergodic theory and introduce fundamental aspects of free probability theory. We explain why asymptotic freeness should be incorporated into the quantum ergodic hierarchy and explore its connection to operator spectral statistics through the predictions of free convolution. In Sections 4, 5, and 6, we analyze the spectral properties of operators in three quantum many-body systems: the mixed-field Ising model with a random magnetic field, a higher-spin variant of the mixed-field Ising model, and the SYK model. Finally, in Section 7, we summarize our results and outline the directions for future research. Technical details and reviews of previous works are deferred to the appendices.
@@ 2.1 Classical ergodic hierarchy
In classical ergodic theory, dynamical systems are characterized by the asymptotic behavior of connected correlation functions at large times. Intuitively, non-integrable dynamics progressively erases correlations between past and future events, rendering them statistically independent as time evolves.
@@
[…]
@@
Classical dynamical systems can be classified based on the degree of chaos they exhibit. This classification is known as the ergodic hierarchy, which organizes systems according to progressively stronger notions of chaotic behavior. For physical systems, the hierarchy includes:
[…]
@@
Although these systems are mathematically rich and provide deep insights into the dynamics of chaos, they are not directly relevant to the discussions in this paper. Thus, we omit a detailed analysis of these classes and focus on the aspects of the ergodic hierarchy most pertinent to our work. For a detailed discussion of the classical ergodic hierarchy, along with a more extensive list of dynamical systems, we refer the reader to Gesteau:2023rrx ; Ouseph:2023juq .
@@ 2.2 Mixing and the Vanishing of Classical Cumulants
as $t\to\infty$ for all $A,B\subset\Sigma$. Similarly, 3-mixing implies the statistical independence of the three events $A$, $B_{t}$, and $C_{t+s}$, which leads to the vanishing of the third classical cumulant:
@@
In the classical case, thinking about mixing in terms of the vanishing of classical cumulants is not strictly necessary. However, this perspective becomes crucial when considering quantum systems, where the concept of statistical independence is replaced by the notion of free independence, which relies on the vanishing of certain mixed free cumulants involving non-crossing partitions.
@@ 2.3 Functions in Phase Space and the Algebraic Perspective
Given a classical probability space $(\Omega,\Sigma,\mu)$, one can define real-valued random variables $f:\Omega\to\mathbb{R}$ as measurable functions that map the sample space to real values. The set of such measurable functions forms an algebra. A notable example is the Banach algebra $L^{\infty}(\Omega,\Sigma,\mu)$ (abbreviated as $L^{\infty}$), consisting of essentially bounded functions. The norm of a function in this algebra is defined as $||f||_{L^{\infty}}=\sup|f|$, and is equipped with a trace functional $\varphi:L^{\infty}\to\mathbb{R}$, given by
@@
Ergodicity and mixing can be discussed naturally in terms of measurable functions. First, a measure-preserving flow satisfies $\int_{x\in\Omega}f(x)\,\textrm{d}\mu=\int_{x\in\Omega}f(T_{t}(x))\,\textrm{d}\mu$. Given $f,g\in L^{\infty}$, one defines the connected correlation function (cumulant) between them as:
@@ 3 Quantum Ergodic Theory
In this section, we present a mathematical framework for quantum dynamical systems, highlighting the fundamental role of non-commutative probability spaces. Additionally, we provide a brief overview of key aspects of free probability theory that are relevant to this work. For a more comprehensive discussion of free probability theory, we refer the reader to speicher2019lecturenotesfreeprobability ; mingo2017free ; nica2006lectures ; Wang:2022ots .
@@
$\lim_{T\rightarrow\infty}\int_{0}^{T}F_{2}(a,b;t)\textrm{d}t=0.$ (13)
@@
Mixing is expected to occur in systems with a continuous spectrum whose algebra is of type III. This is consistent with the fact that in quantum mechanical systems with type I algebras, the spectrum of energies is discrete, and therefore two-point functions are periodic or quasi-periodic functions of time, and cannot decay to zero Berry2001 . Indeed, in the context of modular flows, it can be shown that quantum strong 2-mixing implies that the algebra is of type III Ouseph:2023juq . This is illustrated schematically in Figure 1.
Quantum K-mixing system
A quantum dynamical system $(\mathcal{A},\varphi,T_{t})$ is said to be quantum K-mixing if, for all subalgebras $\mathcal{B}\subset\mathcal{A}$ and all $b\in\mathcal{B}$:
$\lim_{t\rightarrow\infty}\sup_{a\in\mathcal{A}_{(t,\infty)}}|F_{2}(a,b;t)|=0\,,$ (15)
where $\mathcal{A}_{(t,\infty)}=\{T_{s}(A)\mid A\in\mathcal{A},s>t\}$ denotes the time-band algebra associated with the interval $(t,\infty)$. That condition implies that correlations with the entire future decay to zero. Since the K-mixing condition is stronger than 2-mixing, it is likewise expected to occur only for type III algebras.
Since classical K-systems are characterized by the Lyapunov exponent, it is reasonable to expect that quantum K-systems should likewise be characterized by a quantum analogue of the Lyapunov exponent. Indeed, for certain systems, the squared commutator $C_{4}(t)=-\langle[a(0),b(t)]^{2}\rangle$ – motivated by a semiclassical analogy with Poisson brackets larkin1969sov – has been observed to exhibit exponential growth within a specific time window (see for example Kobrin:2020xms ; Craps:2019rbj ). This growth, however, cannot persist indefinitely, as quantum interference effects eventually disrupt the classical picture Richter:2022sik . Notably, this definition of a quantum Lyapunov exponent differs from its classical counterpart Rozenbaum:2016mmv , as it can lead to exponential behavior even in some classically integrable systems, a phenomenon known as saddle-dominated scrambling. This issue can be mitigated by considering logarithmic OTOCs Trunin:2023xmw . In the next subsection, we will discuss how the vanishing of the OTOCs, or equivalently the saturation of the squared commutator $C_{4}(t)=-\langle[a(0),b(t)]^{2}\rangle$ to a constant value, typically associated with chaotic dynamics, can be understood as a consequence of a concept known as asymptotic freeness, a form of statistical independence in quantum systems.
@@
[…]
@@
However, it is important to note that the weakest form of chaos in this hierarchy, namely ergodicity, requires the algebra of observables to be of type II, while stronger notions of chaos necessitate an algebra of type III Ouseph:2023juq . In contrast, finite-dimensional quantum systems exhibit an algebra of type I. As a result, there is no quantum chaos in the sense of mixing for finite-dimensional systems. This is why Michael Berry famously stated, “There is no quantum chaos, only quantum chaology” in his well-known paper berry1989quantum . However, under quantization, classically chaotic systems (specifically K-mixing systems) retain memory of their classical chaotic nature, exhibiting correlations in their spectra that resemble those found in random matrix theory BGS . In the thermodynamic limit, the algebra can change from type I to type II or III Witten:2021jzq , allowing the emergence of ergodicity, mixing, and K-mixing. Algebras of type III also arise in quantum field theory and gauge theories at large $N$ Witten:2021jzq . For such systems, with an infinite number of degrees of freedom, one can truly talk about quantum chaos in a way that is parallel to classical chaos because the large-$N$ limit provides a sort of semiclassical limit Richter:2022sik . See Fig. 2 for a comparison with the classical ergodic hierarchy in relation to number of degrees of freedom and “quantumness” of the classical action.
@@
Figure 2: Illustration of the position of the classical and quantum ergodic hierarchies in relation to the number of degrees of freedom $N$, and classical action $S/\hbar$. The classical ergodic hierarchies deal with classical dynamical systems which in the diagram exist in the classical limit $\hbar/S\rightarrow 0$ (right). In contrast, the quantum ergodic hierarchies deal with non-commuting variables and exist only in the strict limit $1/N\rightarrow 0$ (top). As probes of quantum chaos in quantum systems with $01/N>$ 0, energy-level statistics (arising from the BGS conjecture regarding quantum systems with chaotic classical analogues) and operator spectral statistics do not directly overlap with the quantum ergodic hierarchies. There are different ways in which one can reach an infinite number of degrees of freedom: by considering a lattice model with $L$ sites and taking $L\to\infty$; by taking the continuum limit, where the lattice spacing vanishes, as in QFTs; or by considering a gauge theory matrix/vector model with $N_{c}\to\infty$, where $N_{c}$ is the rank of the gauge group. Note that the parameter $1/N$ plays a role analogous to $\hbar$ in the semi-classical analysis. Diagram inspired by Figure 2 of Richter:2022sik .

### v3

@@ 1 Introduction
Despite its significance across various domains, quantum chaos remains poorly understood. Several diagnostics have been proposed, including spectral statistics governed by random matrix theory BGS , ETH DAlessio:2015qtq , OTOCs Shenker:2014cwa , Krylov complexityParker:2018yvk , and spread complexity Balasubramanian:2022tpr . While these probes offer valuable insights, their interconnections are not well established. [^1] For previous work in this direction, see Cotler:2017jue , which relates averaged OTOCs to the spectral form factor, and rozenbaum2019universal , which discusses the spectral statistics of Lyapunovian operators. Other works in this direction include Ma:2019ocx ; Lau:2022med . These topics are reviewed in Appendices D and F, respectively. It remains unclear whether an overarching theoretical framework exists to unify these different manifestations of quantum chaos or at least provide a systematic classification of them. In many respects, our current understanding of quantum chaos is largely phenomenological, relying on conjectures and heuristic arguments often inspired by semiclassical intuition. For instance, periodic orbit theory has been successfully employed to derive random matrix behavior from classical periodic orbits Berry1985 . However, this approach inherently assumes the presence of well-defined classical trajectories and thus breaks down in systems without a classical limit. Moreover, the mathematically rigorous framework for characterizing chaos – quantum ergodic theory (see Gesteau:2023rrx ; Ouseph:2023juq for recent reviews) – remains largely disconnected from the aforementioned diagnostics of quantum chaos. Developing a mathematically precise and conceptually unified theory of quantum chaos is, therefore, a crucial open challenge, one that could provide a systematic framework for addressing many outstanding questions in the field.
@@
In this work, we propose a quantum ergodic theory based on free probability theory mingo2017free ; voiculescu1992free . This approach is motivated by the fact that any quantum dynamical system can be viewed as a non-commutative probability space equipped with a time evolution map. Since free probability theory studies precisely such non-commutative probability spaces, it is natural to expect its relevance in quantum-mechanical systems. By analogy with classical ergodic theory, chaos can be characterized by how much operators in the future, $B(t)$, become statistically independent of operators in the past, $A(0)$. In free probability theory, a fundamental notion of statistical independence is freeness, meaning that all mixed free cumulants involving the corresponding random variables vanish. In fact, various manifestations of quantum chaos, such as the ETH and the vanishing of OTOCs, can be understood as different avatars of freeness Jindal:2024zcg . Strictly speaking, freeness only emerges in the thermodynamic limit, where the rank $\mathcal{D}$ of the operators in a quantum many-body system tends to infinity. In this context, one speaks of asymptotic freeness, indicating that freeness holds as $\mathcal{D}\to\infty$. At finite $\mathcal{D}$, a hallmark of asymptotic freeness is that the spectrum of combinations of asymptotically free operators follows the predictions of free probability theory Chen:2024zfj . An important first application of free probability theory to quantum chaos was its role in developing generalized versions of the ETH to describe the behavior of OTOCs Pappalardi:2022aaz . This perspective has since been explored in subsequent works PhysRevX.15.011031 ; Vallini:2024bwp ; Pappalardi:2023nsj (see also Wang:2022ots for a discussion of the connection between the gravitational replica trick and free probability).
@@
This work is organized as follows. In Section 2, we review classical ergodic theory. In Section 3 and 4, we discuss quantum ergodic theory and briefly introduce some fundamental aspects of free probability theory that will be used in this paper. We explain why asymptotic freeness should be incorporated into the quantum ergodic hierarchy and explore its connection to operator spectral statistics through the predictions of free convolution. In Sections 5, 6, and 7, we analyze the spectral properties of time-evolved operators in three quantum many-body systems: the mixed-field Ising model with a random magnetic field, a higher-spin variant of the mixed-field Ising model, and the SYK model, and compare the predictions of these statistics obtained from the free probability theory. Finally, in Section 8, we summarize our results and outline the directions for future research. Some technical details and reviews of previous works that are relevant to the present work are deferred to the appendices.
@@ 2.1 Classical ergodic hierarchy
In classical ergodic theory, dynamical systems are characterized by the asymptotic behavior of connected correlation functions at large times. Intuitively, non-integrable dynamics progressively erases correlations between past and future events, rendering them statistically independent as time evolves berkovitz2006ergodic .
@@
The definition of ergodicity given in (2) is equivalent to the following definition: a classical dynamical system is ergodic if, for every $T_{t}$-invariant set $A\subset\Sigma$ (i.e. $T_{t}(A)=A$) has either $\mu(A)=0$ or $\mu(A)=1$ [^2] See Section 1.6 of walters1982introduction for an extensive discussion of equivalent notions of ergodicity.. For an ergodic system, the only subsets of $\Sigma$ that are invariant under $T_{t}$ are the whole phase space $\Omega$ and the sets of zero measure. This implies that a typical trajectory cannot be confined to a restricted region of the phase space but instead explores the entire phase space, except for a subset of measure zero. [^3] As explained in bohigas1984chaotic , the strict form of Boltzmann’s ergodic hypothesis can never hold, since a trajectory that passes through every point in $\Omega$ would intersect itself, which is impossible. In particular, this implies the equality between time averages and phase space averages. We will discuss this in more detail in Section 2.3 after introducing functions in phase space.
@@
Classical dynamical systems can be classified based on their mixing properties. This classification is known as the ergodic hierarchy, which organizes systems according to progressively stronger notions of mixing behavior. For physical systems, the hierarchy includes:
$\text{K-mixing systems}\subset\text{Strongly mixing systems}\subset\text{Ergodic systems}.$
@@
Although these systems are mathematically rich and provide deep insights into the dynamics of chaos, they are not directly relevant to the discussions in this paper. Thus, we omit a detailed analysis of these classes and focus on the aspects of the ergodic hierarchy most pertinent to our work. For a detailed discussion of the classical ergodic hierarchy, its connection with randomness, along with a more extensive list of dynamical systems, we refer the reader to berkovitz2006ergodic ; Gesteau:2023rrx ; Ouseph:2023juq .
@@ 2.2 Mixing and the Vanishing of Classical Cumulants
as $t\to\infty$ for all $A,B\subset\Sigma$. Similarly, 3-mixing (along with 2-mixing) implies the statistical independence of the three events $A$, $B_{t}$, and $C_{t+s}$, which leads to the vanishing of the third classical cumulant:
@@
In the classical case, thinking about mixing in terms of the vanishing of classical cumulants is not strictly necessary. However, as we argue in the following, this perspective becomes crucial when considering quantum systems, where the concept of statistical independence is replaced by the notion of free independence, which relies on the vanishing of certain mixed free cumulants involving non-crossing partitions.
@@ 2.3 Functions in Phase Space and the Algebraic Perspective
Given a classical probability space $(\Omega,\Sigma,\mu)$, one can define real-valued random variables $f:\Omega\to\mathbb{R}$ as measurable functions that map the sample space to real values. The set of such measurable functions forms an algebra. A notable example is the Banach algebra $L^{\infty}(\Omega,\Sigma,\mu)$ (abbreviated as $L^{\infty}$), consisting of essentially bounded functions. The norm of a function in this algebra is defined as $f||_{L^{\infty}}=\sup|f|$, and is equipped with a trace functional $\varphi:L^{\infty}\to\mathbb{R}$, given by
@@
Ergodicity and mixing can be discussed naturally in terms of measurable functions. First, a measure-preserving flow satisfies $\int_{x\in\Omega}f(x)\,\textrm{d}\mu=\int_{x\in\Omega}f(T_{t}(x))\,\textrm{d}\mu$. Given $f,g\in L^{\infty}$, one can define the connected correlation function (cumulant) between them. E.g., the second-order cumulant can be written as:
@@ 3 Quantum Ergodic Theory
In this section, we present a mathematical framework for quantum dynamical systems, highlighting the fundamental role of non-commutative probability spaces. Additionally, we provide a brief overview of key aspects of free probability theory that are relevant to this work. For a more comprehensive discussion of free probability theory, we refer the reader to speicher2019lecture ; mingo2017free ; nica2006lectures ; voiculescu1992free .
@@
$\lim_{T\rightarrow\infty}\int_{0}^{T}F_{2}(a,b;t)~\textrm{d}t=0.$ (13)
@@
Mixing is expected to occur in systems with a continuous spectrum whose algebra is type II or III. This is consistent with the fact that in quantum mechanical systems with type I algebras, the spectrum of energies is discrete, and therefore two-point functions are periodic or quasi-periodic functions of time, and cannot decay to zero Berry2001 . Indeed, in the context of modular flows, it can be shown that quantum strong 2-mixing implies that the algebra is of type III Ouseph:2023juq . This is illustrated schematically in Figure 1.
Algebraic Quantum K-system Narnhofer:2001
A quantum dynamical system $(\mathcal{A},\varphi,T_{t})$ with a time-invariant state $\varphi$ is called an algebraic quantum K-system if there exists a subalgebra $\mathcal{A}_{0}\subset\mathcal{A}$ such that, for all $t\geq 0$:
$T_{t}(\mathcal{A}_{0})\supset\mathcal{A}_{0},\quad\bigcup_{t\geq 0}T_{t}(\mathcal{A}_{0})=\mathcal{A},\quad\text{and}\quad\bigcap_{t\geq 0}T_{-t}(\mathcal{A}_{0})=\mathbb{C}\,\mathbb{I}.$ (15)
Here, $\mathbb{C}\,\mathbb{I}$ denotes the scalar multiples of the identity. Quantum K-systems exhibit quantum $n$-mixing (see Theorem 2.3 in Narnhofer:2001 ), meaning that for any finite collection of operators $A_{1},A_{2},\ldots,A_{n}\in\mathcal{A}$, the following holds :
$\lim_{\begin{subarray}{c}t_{i+1}-t_{i}\to\infty\\ i=1,\ldots,n-1\end{subarray}}\varphi\big{(}A_{1}(t_{1})A_{2}(t_{2})\cdots A_{n}(t_{n})\big{)}=\prod_{i=1}^{n}\varphi(A_{i}),$ (16)
where $A_{i}(t):=T_{t}(A_{i})$.
Since classical K-systems are typically characterized by Lyapunov exponents, it is reasonable to expect that quantum K-systems should likewise be characterized by a quantum analogue of the Lyapunov exponent. Indeed, for certain systems, the squared commutator $C_{4}(t)=-\langle[a(0),b(t)]^{2}\rangle$ – motivated by a semiclassical analogy with Poisson brackets larkin1969sov – has been observed to exhibit exponential growth within a specific time window (see for example Kobrin:2020xms ; Craps:2019rbj ). This growth, however, cannot persist indefinitely, as quantum interference effects eventually disrupt the classical picture Richter:2022sik . Notably, this definition of a quantum Lyapunov exponent differs from its classical counterpart Rozenbaum:2016mmv , as it can lead to exponential behavior even in some classically integrable systems, a phenomenon known as saddle-dominated scrambling. This issue can be mitigated by considering logarithmic OTOCs Trunin:2023xmw . In the next subsection, we will discuss how the vanishing of the OTOCs, or equivalently the saturation of the squared commutator $C_{4}(t)=-\langle[a(0),b(t)]^{2}\rangle$ to a constant value, typically associated with chaotic dynamics, can be understood as a consequence of a concept known as asymptotic freeness, a form of statistical independence in quantum systems.
@@
$\text{Quantum K-systems}\subset\text{Quantum Mixing systems}\subset\text{Quantum Ergodic systems}.$
@@
However, it is important to note that the weakest form of chaos in this hierarchy, namely ergodicity, requires the algebra of observables to be of type II, while stronger notions of chaos may necessitate an algebra of type III Ouseph:2023juq . In contrast, finite-dimensional quantum systems exhibit an algebra of type I. As a result, there is no quantum chaos in the sense of mixing for finite-dimensional systems. This is why Michael Berry famously stated, “There is no quantum chaos, only quantum chaology” in his well-known paper berry1989quantum . However, under quantization, classically chaotic systems (specifically K-mixing systems) retain memory of their classical chaotic nature, exhibiting correlations in their spectra that resemble those found in random matrix theory BGS . In the thermodynamic limit, the algebra can change from type I to type II or III Witten:2021jzq , allowing the emergence of quantum ergodicity and mixing. Algebras of type III also arise in quantum field theory and gauge theories at large $N$ Witten:2021jzq . For such systems, with an infinite number of degrees of freedom, one can truly talk about quantum chaos in a way that is parallel to classical chaos because the large-$N$ limit provides a sort of semiclassical limit Richter:2022sik . See Fig. 2 for a comparison with the classical ergodic hierarchy in relation to number of degrees of freedom and “quantumness” of the classical action. [^7] For a collection of works that appeared in the physics and mathematics literature on the types of systems that appear in the quantum ergodic hierarchy and related to the discussion presented above, see narnhofer1989mixing ; narnhofer1989quantum ; Narnhofer:2001 ; zaslavsky1981stochasticity ; peres1984ergodicity ; castagnino2009towards ; vikram2023dynamical ; benatti1998statistics . Also, for recent progress in understanding quantum ergodic hierarchy in the context of unitary quantum circuits, a solvable model for many-body quantum chaos, we refer to Aravinda:2021sxc ; Aravinda:2023vsg .
@@
Figure 2: Illustration of the position of the classical and quantum ergodic hierarchies in relation to the number of degrees of freedom NN, and classical action $S/\hbar$. The classical ergodic hierarchy deals with classical dynamical systems, which in the diagram exist in the classical limit $\hbar/S\rightarrow 0$ (right). In contrast, the quantum ergodic hierarchy deals with non-commuting variables and exists only in the strict thermodynamic/continuum limit $1/N\rightarrow 0$ (top). As probes of quantum chaos in quantum systems with $1/N>0$, energy-level statistics (arising from the BGS conjecture regarding quantum systems with chaotic classical analogues) and operator spectral statistics do not directly overlap with the quantum ergodic hierarchy. There are different ways in which one can reach an infinite number of degrees of freedom: by considering a lattice model with $L$ sites and taking $L\to\infty$; by taking the continuum limit, where the lattice spacing vanishes, as in QFTs; or by considering a gauge theory matrix/vector model with $N_{c}\to\infty$, where $N_{c}$ is the rank of the gauge group. Note that the parameter $1/N$ plays a role analogous to $\hbar$ in the semi-classical analysis. Diagram inspired by Figure 2 of Richter:2022sik .